Pedrodog10:

The first sunrise in the United States each day is observed from Cadillac Mountain on Mount Desert Island in Maine. The time of the sunrise can be modeled by t(m) = 1.665 sin (m + 3) + 5.485, where t is hours after midnight and m is the number of months after January 1. When does the sunrise at 6:00 A.M.?

Creed:

First we will subsitute 6 for variable "t", Then solve for variable "m" using Algebra; \[6 = 1.665\sin \frac{ \Pi }{ 6 }(m+ 3)+ 5.485\] Now, you need to isolate the sine term; \[\frac{ 6-5.485 }{ 1.665 }=\sin \frac{ \Pi }{ 6 }(m+3)\] Now, Imply the Inverse Sine \[\sin^{-1} (0.3093)=\frac{ \Pi }{ 6 }(m+3)\] Sine is positive in Quadrants I and II. \[\sin^{-1} (0.3039)=\frac{ \Pi }{ 6 }(m+3)\] \[0.314=\frac{ \Pi }{ 6} (m+ 3)\] \[(\frac{ 6 }{ \Pi})0.314=m+3\] \[-2.400=m\] & \[2.400=m\]

Creed:

𝘐 𝘩𝘰𝘱𝘦 𝘵𝘩𝘢𝘵 𝘩𝘦𝘭𝘱𝘦𝘥, 𝘭𝘦𝘵 𝘮𝘦 𝘬𝘯𝘰𝘸 𝘪𝘧 𝘺𝘰𝘶 𝘯𝘦𝘦𝘥 𝘢𝘯𝘺𝘵𝘩𝘪𝘯𝘨 𝘦𝘭𝘴𝘦 𝘣𝘺 𝘤𝘰𝘯𝘵𝘢𝘤𝘵𝘪𝘯𝘨 𝘮𝘦. ℂ𝕣𝕖𝕖𝕕